OpenStudy (anonymous):

12x^2y^3-6xy ________________(division line) 3xy

OpenStudy (accessdenied):

We can (conveniently) factor out a 3xy from the numerator. 3xy(4xy^2 - 2) ------------- 3xy From there, you can just cancel out the 3xy in numerator and denominator. :)

OpenStudy (anonymous):

could u explain a little more please..

OpenStudy (anonymous):

but how u get 4xy^2

OpenStudy (accessdenied):

When we factor out a 3xy from 12x^2y^3, we would divide it out. \[ \frac{12x^{2}y^{3}}{3xy} \] 12/3 = 4. x^2 / x = x/1. y^3/y = y^2

OpenStudy (anonymous):

oh..i get it...thanks!!!!!!